OpenStudy (anonymous):

Find the square. (2x – 6)2

Parth (parthkohli):

Use the identity: \(\Large \color{MidnightBlue}{\Rightarrow (a - b)^2 = a^2 \pm 2ab + b^2 }\)

Parth (parthkohli):

I mean - 2ab

OpenStudy (eyust707):

\[(2x-6)^2 = (2x -6) (2x -6)\]

OpenStudy (anonymous):

4x2 – 24x + 36? is that right?

OpenStudy (eyust707):

what parth posted is a shortcut if you have the difference of a perfect square

OpenStudy (eyust707):

correct!

OpenStudy (anonymous):

thank you

OpenStudy (eyust707):

np